OpenStudy (anonymous):

Solve for x: (x+20)/2 = = 3x

OpenStudy (anonymous):

Solve for x: (x+20)/2 =3x

OpenStudy (anonymous):

Get all the x's on one side. For a hint, I would start by getting rid of the fraction by eliminating the 2.

OpenStudy (anonymous):

then what

OpenStudy (anonymous):

It's pretty straightforward from there. What did you get when you got rid of the 2?

OpenStudy (anonymous):

i didnt do it yet i just want somebody to give me the step on hoe to do these prop

OpenStudy (anonymous):

\[\frac{ x+20 }{ 2 }=3x \] \[\frac{ x }{ 2 } + \frac{20}{2}=3x \] \[\frac{ x }{ 2 } + 10=3x \] \[-\frac{x}{2}\] to both sides gives us \[10= \frac{ 5x }{ 2 } \] multiply both sides by 2 so 20=5x then 5 divided to both sides gives us 4

OpenStudy (anonymous):

x=4